OpenStudy (anonymous):

evaluate the limit of x as x approaches to 8 is equal to ^(7/2)

OpenStudy (jdoe0001):

^(7/2)?

OpenStudy (anonymous):

lim x^(7/2) as x->8

OpenStudy (jdoe0001):

\(\bf \large {lim_{x\rightarrow 8} x^{\frac{7}{2}}\\ x^{\frac{7}{2}} \implies \sqrt[2]{x^7}\implies \square?}\)

OpenStudy (jdoe0001):

hmm \(\bf \large { lim_{x\rightarrow 8} x^{\frac{7}{3}}\\ x^{\frac{7}{2}} \implies \sqrt[3]{x^7}\implies \square?}\)

OpenStudy (jdoe0001):

woops, missed one 3, well \(\bf\large { lim_{x\rightarrow 8} x^{\frac{7}{3}}\\ x^{\frac{7}{3}} \implies \sqrt[3]{x^7}\implies \square?}\)

OpenStudy (jdoe0001):

keep in mind that \(\bf x^7 \implies x^3 \times x^3 \times x\)

OpenStudy (anonymous):

the system saids it's still wrong

OpenStudy (jdoe0001):

is wrong? we don't even know what's wrong :S

OpenStudy (anonymous):

lol

OpenStudy (jdoe0001):

what did you use?///////////////////////////

OpenStudy (jdoe0001):

darn key :(

OpenStudy (goformit100):

First of all @BaoZhu : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct